An object of 4.50 kg fell for a total of 3.40 seconds before hitting the ground. If the initial velocity of the object was zero, from what height did it fall? (Ignore air resistance)
1
Expert's answer
2020-10-13T09:53:30-0400

Equation of motion of an object is "y(t) = h - \frac{g t^2}{2}", where "h" is the initial height. After time "t = 3.4 s" "y(t)" is zero, hence: "0 = h - \frac{g t^2}{2}", from where "h = \frac{g t^2}{2} \approx 56.7 m".